# Operational research approaches and mathematical models for kidney exchange: A literature survey and empirical evaluation

### v1

@@ 1 Introduction
Whilst the largest possible set of transplants could contain cycles of arbitrary length, longer cycles can be more vulnerable to failure, as it may take only one recipient falling ill, or one unexpected positive crossmatch, to result in a cycle not proceeding, with all associated transplant opportunities potentially being lost. Moreover, to avoid a scenario where a donor donates a kidney while their paired recipient does not receive a kidney due to a failed transplant (or even due to a donor reneging), many KEPs aim to perform all nephrectomies and all transplants associated with a given cycle simultaneously. Longer cycles can thus create difficult logistical challenges requiring the simultaneous scheduling of many distinct surgeries. For these reasons, many KEPs apply a strict upper limit on the number of transplants that are included in any one cycle (Biró et al., , 2021). (Note that if cycles do fail, re-optimisation is one strategy, whilst recourse, described further below, is another (Pedroso, , 2014).)
@@
The transplants associated with a chain can be performed non-simultaneously, such that each recipient receives a kidney donation before their paired donor donates a kidney. This means there is less risk associated with longer chains compared to longer cycles, so KEPs can sometimes allow chains to be longer than cycles (Biró et al., , 2021). We refer to an exchange as a cycle or chain in a KEP. Note that a recipient in a KEP may have multiple willing but incompatible donors. Henceforth we assume that each recipient only has one paired donor. We explain in Section 3.1 why this is without loss of generality. Moreover sometimes medically compatible RDPs participate in a KEP (Gentry et al., , 2007), with the aim of the recipient obtaining a better match (e.g., from a younger donor) than from their paired donor – such pairs can also help to construct exchanges involving incompatible RDPs that would not otherwise exist.
Rapaport, (1986) first introduced the concept of kidney exchange, also sometimes referred to in the literature as kidney paired donation, suggesting the possibility of pairwise exchanges, which are cycles involving two RDPs. The first kidney exchanges in the world were carried out in South Korea in 1991 (Kwak et al., , 1999), whilst the first European kidney exchange occurred in Switzerland in 1999 (Thiel et al., , 2001), and the first US kidney exchange was carried out the following year (Zarsadias, , 2010). The first national KEP was established in the Netherlands in 2004 (de Klerk et al., , 2005). Establishing a KEP normally requires ethical and legal hurdles to be overcome (van Basshuysen, , 2020). This is because a kidney exchange will typically involve a recipient obtaining a kidney from a donor who is not known to them, whereas it is usually the case that a direct donation from a living donor to a recipient is only possible where the recipient has an emotional attachment or a genetic connection to the donor (e.g., they are a spouse or blood relative). By way of example, the UK’s KEP began in 2007 following the introduction of the Human Tissue Act (2004) and Human Tissue (Scotland) Act (2006), which provided the legal framework to enable transplants between strangers in the absence of financial reward.
@@
KE-Opt is often studied using a graph-theoretic model involving the underlying compatibility graph $\mathcal{G}$, which contains a vertex for each NDD and RDP, and an arc $(u,v)$ from an NDD or RDP $u$ to an RDP $v$ whenever the donor of $u$ is medically compatible with the recipient of $v$. A set of (kidney) exchanges is then a vertex-disjoint set of cycles and chains in $\mathcal{G}$. In some KEP settings, the arcs of the compatibility graph have weights associated with them, usually representing the utilities of the associated potential transplants. The first papers to study algorithms or mechanisms for KE-Opt were the landmark papers of Roth et al., (2004, 2005). When the objective is to maximise the number of transplants, KE-Opt is $\mathcal{NP}$-hard for a fixed upper bound of 3 or more on the cycle length, even if there are no NDDs (Abraham et al., , 2007).
@@
  • A detailed literature survey (with over 190 references) of OR approaches to KE-Opt, covering the following topics: algorithms and complexity for KE-Opt; hierarchical optimisation in KE-Opt; enabling equal access to transplantation; dynamic KEPs; robustness in KEPs; multi-hospital and international KEPs; recipients’ preferences; dataset generators and software tools; emerging topics; and other related surveys.
  • A systematic exposition of all the key existing ILP approaches for KE-Opt, describing separately models for representing optimal solutions comprising only cycles from those comprising only chains. As a consequence, combined ILP models for KE-Opt can be obtained by mixing a cycle model with a chain model. We also use a running example to illustrate all models for the benefit of the reader.
  • A comprehensive empirical evaluation of all combined ILP models for KE-Opt that are described in this paper, together with “off-the-shelf” approaches involving advanced techniques such as column generation and branch-and-price, where we have been able to obtain and execute the third-party software. The main aim is to compare execution times of the different approaches considered on randomly generated datasets that reflect the characteristics of real data from the UK KEP. In particular, we tested 49 methods on 4,320 instances, corresponding to over 200,000 algorithm executions, and amounting to over 10 years of computational processing time.
@@
The remainder of this paper is structured as follows. Section 2 contains the literature survey. Section 3 presents a comprehensive exposition of ILP formulations for KE-Opt, separating cycle and chain models and showing how they can be combined. Section 4 presents the empirical evaluation of ILP formulations and third-party methods for KE-Opt, and Section 5 concludes with some directions for future research.
@@ 2.1 Algorithms and Complexity for KE-Opt
OR and computer science play a key role in KEPs, as efficient and effective algorithms to compute sets of exchanges must be designed, taking into account the KEP constraints and optimality criteria. First, we give theoretical results relating to the complexity and approximability of KE-Opt. Second, we give an overview of constraint programming and heuristic algorithms for KE-Opt that have been studied in the literature. Note that in this section, we are describing results relating to the version of KE-Opt where we want to find a set of exchanges with the optimal number of transplants given some cycle and chain limits.
@@
Once the connection between KEPs and market clearing was established by Roth et al., (2004), computational complexity theory became a key tool for analysing KEPs. The standard form of KE-Opt takes as input a compatibility graph together with cycle and chain length limits, and as output, we seek a set of exchanges with the maximum number of transplants whilst respecting these cycle and chain length restrictions. Roth et al., (2005) showed that this variant of KE-Opt is solvable in polynomial time when only cycles containing two RDPs are permitted, by reducing to a maximum matching problem in a general graph (see also Gentry et al., (2020)). It is not difficult to extend this result to the case where chains with one RDP are also permitted. Abraham et al., (2007) showed that the problem becomes $\mathcal{NP}$-hard when cycles containing at most $K$ RDPs are permitted, for any $K\geq 3$, even if there are no NDDs. In the case that cycles and chains can have any number of RDPs, this variant of KE-Opt becomes solvable in polynomial time again, by reducing to a maximum matching problem in a bipartite graph (Abraham et al., , 2007). The problem is also solvable in polynomial time when the compatibility graph is represented using a constant number of recipient and donor attributes (Dickerson et al., , 2017).
@@
Xiao and Wang, (2018) provided an exact algorithm to find an optimal set of exchanges, with a running time of $O(2^{n}n^{3})$, where $n$ is the number of RDPs. Further approaches to designing exact algorithms are based on parameterised complexity. Lin et al., (2019) developed two randomised fixed-parameter tractable ($\mathcal{FPT}$) algorithms for KE-Opt without chains; the first is for cycles of length at most $3$ and is parameterised by the number of $2$-cycles and $3$-cycles, whilst the second is for cycles of length at most $K$ (for fixed $K\geq 3$) and is parameterised by the number of transplants. Maiti and Dey, (2022) described $\mathcal{FPT}$ algorithms when parameterised by the number of transplants, vertex types, or the sum of the graph’s treewidth and the maximum cycle or chain limit. Hébert-Johnson et al., (2024) showed that the problem is $\mathcal{W}[1]$-hard when parameterised by only treewidth but $\mathcal{FPT}$ when parameterised by vertex type. They also gave an improved $\mathcal{FPT}$ algorithm when the parameter is the number of transplants.
@@
Another approach is approximation algorithms. Biro et al., (2009) showed that the problem of finding a maximum size set of exchanges with cycles of length at most $3$ is $\mathcal{APX}$-complete, i.e., the optimal solution cannot be approximated within some constant factor. Luo et al., (2016) established inapproximability results for KE-Opt in both the weighted (i.e., we seek a maximum weight set of exchanges in the presence of arc weights in the compatibility graph) and unweighted cases, giving specific lower bounds beyond which KE-Opt is not approximable unless $\mathcal{P}=\mathcal{NP}$. Jia et al., (2017) provided a black-box reduction linking the cycle packing problem (a reformulation of the problem of finding a maximum size set of exchanges in a KEP instance) to set packing, leading to a $(3/2+\varepsilon)$-approximation algorithm for cycle limit $3$ and a $(7/3+\varepsilon)$-approximation when there is a cycle length limit of $4$.
@@ 2.2 Hierarchical Optimisation in KE-Opt
For many countries, their KEPs do not solely aim to maximise the number of transplants found in a given matching run. Instead, they typically find a set of exchanges that is optimal with respect to several objectives that take into account the number of transplants selected, the solution’s structure, and notions of fairness and recourse. OR provides a range of methods to optimise over multiple objectives, and many KEPs use hierarchical optimisation (Biró et al., , 2021). This type of multi-objective optimisation sequentially optimises the objectives according to some priority order; when optimising a given objective, the optimal values for the previous objectives are maintained.
@@
The hierarchical objectives used in KEPs and their orderings can differ greatly between countries; the survey by Biró et al., (2021) highlighted these differences across Europe. For example, the Spanish national KEP uses four objectives and optimises them hierarchically. Given the pool of RDPs, they first find the maximum number of transplants, say $n_{T}$. They then find the maximum number $n_{E}$ of distinct exchanges in a solution with $n_{T}$ transplants (forcing selected chains and cycles to be shorter). Following this, they find a solution with $n_{T}$ transplants, $n_{E}$ distinct exchanges, and a maximum number of cross-arcs. [^1] A cross-arc is an additional arc among the NDD/RDPs in an exchange $c$, but not already an arc of $c$, that allows an alternative exchange $c^{\prime}$ to take place in the case of a vertex or arc failure in $c$ (e.g., if a donor becomes ill, or if a laboratory crossmatch is identified). Including an objective that maximises the number of cross-arcs in a solution gives additional recourse possibilities if some transplants cannot proceed. Finally, while maintaining all the previous optimal values, the chosen solution has an optimal total weight with respect to weights associated with the selected arcs.
@@
Given that hierarchical optimisation often leads to multiple $\mathcal{NP}$-hard problems being solved in succession, various OR techniques have been employed to find optimal solutions. Manlove and O’Malley, (2014) extended the classical cycle formulation given by Roth et al., (2007) to be able to find solutions according to the UK’s optimality criteria. They then analysed the impact of extending the current criteria to allow cycles of length $4$ instead of $3$ and showed that this could increase the total number of transplants on real data from the UK’s national KEP. Delorme et al., (2024) developed four strategies to reduce the running time of ILP-based algorithms when dealing with hierarchical optimisation: eliminating the dominated exchanges (cycles or chains that cannot appear in any optimal solution given the considered criteria), objective diving (using dual bounds to set the value of early objectives, possibly backtracking if a fixed value is, in fact, infeasible), reduced-cost variable fixing (see Section 3.3), and model swapping (using a different ILP model depending on the objective currently optimized). The effectiveness of their approach was demonstrated on three sets of hierarchical objective functions: from the UK, Spain, and the Netherlands. A similar line of research was conducted by Glorie et al., 2014b , who created an iterative branch-and-price algorithm to find optimal exchanges for the Dutch national KEP criteria (see also Section 3.3). Klimentova et al., (2014) also studied hierarchical optimisation that first maximises the number of transplants and then maximises the number of exchanges (a proxy for minimising the average exchange length). They solved this hierarchical optimisation problem by creating a single combined objective function with a smart choice of weights so that the objective functions are solved hierarchically, using a branch-and-price algorithm. Interestingly, their results seem to indicate that optimising on a combination of the two objectives simultaneously (using appropriate coefficients in the objective function) leads to better results than solving sequentially.
@@ 2.3 Enabling Equal Access to Transplantation
The next area of the literature we explore relates to ensuring that recipients have equal access to transplantation, and thus, many of the cited works provide an economic perspective of KEPs. There can be many reasons why some recipients could be less likely to be selected for a transplant in a KEP. So-called hard-to-match recipients tend to have fewer compatible donors in the pool. When only maximising the number of transplants in a matching run, hard-to-match recipients could repeatedly not be selected, raising concerns about fairness for these recipients.
@@
One way to enable equal access to transplantation involves altering the mechanism to prioritise the likelihood of certain recipients being selected. Roth et al., (2005) were the first to study algorithmic fairness in KEPs. Their egalitarian mechanism maximises the potential utility of harder-to-match RDPs when KEPs allow for stochastic outcomes, i.e. when probabilities are given over possible outcomes. Li et al., (2014) provided polynomial time algorithms to combat the exponential algorithm given by Roth et al., (2005), while keeping a level of equity among the RDPs in their likelihood of being selected. More recently, Demeulemeester et al., (2025) studied fairness in KEPs when there are multiple optimal solutions and the solution is selected by a general-purpose mathematical solver. They demonstrated that not all optimal solutions had the same probability of being returned by the solver and proposed several algorithms to alleviate this issue. A similar approach was taken by Farnadi et al., (2021), whose focus was on the equity of access to transplantation, especially in the presence of multiple optimal solutions. They enumerated all optimal solutions by using a hybrid of constraint programming and linear programming to avoid any bias that a deterministic solver may have.
@@
Highly sensitised recipients have a low level of tissue-type compatibility and therefore, tend to be harder to match. As such, studies have focused on mechanisms that relax the objective to maximise the number of transplants, in order to help these recipients. Dickerson et al., (2014) were the first to introduce the price of fairness (PoF) to KEPs, a measure that quantifies the trade-off between efficiency (maximising the number of transplants) and fairness (improving access to transplantation for highly-sensitised recipients). McElfresh and Dickerson, (2018) extended upon this by studying the PoF when there are also NDDs, showing that the PoF tends to zero when there are many NDDs. On the other hand, Ashlagi et al., (2012) showed through theoretical results and via simulations that longer chains could help select more highly sensitised recipients without negatively impacting the likelihood of the other RDPs from being selected. However, McElfresh and Dickerson, (2018) showed that the mechanisms by Dickerson et al., (2014) have an arbitrarily bad PoF when the length of cycles and chains increase. Yet, their findings suggest that fairer solutions for highly sensitised patients tend to use longer chains and cycles. The mechanisms created by McElfresh and Dickerson, (2018) were shown to limit the loss in efficiency in terms of PoF when directly prioritising disadvantaged recipients. Duppala et al., (2023) presented randomised polynomial-time algorithms that take into account proportionality vectors that indicate how to prioritise certain groups, returning a probabilistically fair solution with provable guarantees.

### v2

@@ 1 Introduction
Whilst the largest possible set of transplants could contain cycles of arbitrary length, longer cycles can be more vulnerable to failure, as it may take only one recipient falling ill, or one unexpected positive crossmatch (which is an indicator of tissue-type incompatibility between the donor and recipient), to result in a cycle not proceeding, with all associated transplant opportunities potentially being lost. Moreover, to avoid a scenario where a donor donates a kidney while their paired recipient does not receive a kidney due to a failed transplant (or even due to a donor reneging), many KEPs aim to perform all nephrectomies and all transplants associated with a given cycle simultaneously. Longer cycles can thus create difficult logistical challenges requiring the simultaneous scheduling of many distinct surgeries. For these reasons, many KEPs apply a strict upper limit on the number of transplants that are included in any one cycle (Biró et al., , 2021). (Note that if cycles do fail, re-optimisation is one strategy, whilst recourse, described further below, is another (Pedroso, , 2014).)
@@
The transplants associated with a chain can be performed non-simultaneously, such that each recipient receives a kidney donation before their paired donor donates a kidney. This means there is less risk associated with longer chains compared to longer cycles, so KEPs can sometimes allow chains to be longer than cycles (Biró et al., , 2021). We refer to an exchange as a cycle or chain in a KEP. Note that a recipient in a KEP may have multiple willing but incompatible donors. Henceforth we assume that each recipient only has one paired donor , and we explain in Section 3.1 why , in most cases, we lose no generality by making this assumption. Moreover sometimes medically compatible RDPs participate in a KEP (Gentry et al., , 2007), with the aim of the recipient obtaining a better match (e.g., from a younger donor) than from their paired donor – such pairs can also help to construct exchanges involving incompatible RDPs that would not otherwise exist.
Rapaport, (1986) first introduced the concept of kidney exchange, also sometimes referred to in the literature as kidney paired donation, suggesting the possibility of pairwise exchanges, which are cycles involving two RDPs. The first kidney exchanges in the world were carried out in South Korea in 1991 (Kwak et al., , 1999), whilst the first European kidney exchange occurred in Switzerland in 1999 (Thiel et al., , 2001), and the first US kidney exchange was carried out the following year (Zarsadias, , 2010). The first national KEP was established in the Netherlands in 2004 (de Klerk et al., , 2005). Establishing a KEP normally requires ethical and legal hurdles to be overcome  (Ross et al., , 1997; Ross and Woodle, , 2000; van Basshuysen, , 2020). This is because a kidney exchange will typically involve a recipient obtaining a kidney from a donor who is not known to them, whereas it is usually the case that a direct donation from a living donor to a recipient is only possible where the recipient has an emotional attachment or a genetic connection to the donor (e.g., they are a spouse or blood relative). By way of example, the UK KEP began in 2007 following the introduction of the Human Tissue Act (2004) and Human Tissue (Scotland) Act (2006), which provided the legal framework to enable transplants between strangers in the absence of financial reward.
@@
KE-Opt is often studied using a graph-theoretic model involving the underlying compatibility graph $\mathcal{G}$, which contains a vertex for each NDD and RDP, and an arc $(u,v)$ from an NDD or RDP $u$ to an RDP $v$ whenever the donor of $u$ is medically compatible with the recipient of $v$. A set of (kidney) exchanges is then a vertex-disjoint set of cycles and chains in $\mathcal{G}$. In some KEP settings, the arcs of the compatibility graph have weights associated with them, usually representing the utilities of the associated potential transplants. The first papers to study algorithms or mechanisms for KE-Opt were the landmark papers of Roth et al., (2004, 2005). When the objective is to maximise the number of transplants, KE-Opt is $\mathcal{NP}$-hard in general (Abraham et al., , 2007).
@@
  • A detailed literature survey (with over 210 references) of OR approaches to KE-Opt, covering the following topics: algorithms and complexity for KE-Opt; hierarchical optimisation in KE-Opt; enabling equal access to transplantation; dynamic KEPs; uncertainty and robustness in KEPs; multi-hospital and international KEPs; recipients’ preferences; dataset generators and software tools; emerging topics; and other related surveys.
  • A systematic exposition of all the key existing ILP approaches for KE-Opt, describing separately models for representing optimal solutions comprising only cycles from those comprising only chains. As a consequence, combined ILP models for KE-Opt can be obtained by mixing a cycle model with a chain model. We also use a running example ( appearing in Appendix D) to illustrate all models for the benefit of the reader.
  • A comprehensive empirical evaluation of all combined ILP models for KE-Opt that are described in this paper, together with “off-the-shelf” approaches involving advanced techniques such as column generation and branch-and-price, where we have been able to obtain and execute the third-party software. The main aim is to compare execution times of the different approaches considered on randomly generated datasets that reflect the characteristics of real data from the UK’s KEP. In particular, we tested 49 methods on 4,320 instances, corresponding to over 200,000 algorithm executions, and amounting to over 10 years of computational processing time in total, across multiple cores running in parallel.
@@
The remainder of this paper is structured as follows. Section 2 contains the literature survey. Section 3 presents a comprehensive exposition of ILP formulations for KE-Opt, separating cycle and chain models and showing how they can be combined. Section 4 presents the empirical evaluation of ILP formulations and third-party methods for KE-Opt, and Section 5 concludes with some directions for future research. For space reasons, our running example, some technical exposition and part of the experimental results are contained in the appendices.
@@ 2.1 Algorithms and Complexity for KE-Opt
OR and computer science play a key role in KEPs, as efficient and effective algorithms to compute sets of exchanges must be designed, taking into account the KEP constraints and optimality criteria. First, we give theoretical results relating to the complexity and approximability of KE-Opt. Second, we give an overview of constraint programming and heuristic algorithms for KE-Opt that have been studied in the literature. Note that in this section, we are describing results relating to the version of KE-Opt where we seek a set of exchanges with the maximum number of transplants , given some cycle and chain length limits.
@@
Once the connection between KEPs and market clearing was established by Roth et al., (2004), computational complexity theory became a key tool for analysing KEPs. The standard form of KE-Opt takes as input a compatibility graph together with cycle and chain length limits, and as output, we seek a set of exchanges with the maximum number of transplants whilst respecting these cycle and chain length restrictions. Roth et al., (2005) showed that this variant of KE-Opt is solvable in polynomial time when only cycles containing two RDPs are permitted, by reducing to a maximum matching problem in a general graph (see also Gentry et al., (2020)). It is not difficult to extend this result to the case where chains with one RDP are also permitted. Abraham et al., (2007) showed that the problem becomes $\mathcal{NP}$-hard when cycles containing at most $K$ RDPs are permitted, for any $K\geq 3$, even if there are no NDDs. In the case that cycles and chains can have any number of RDPs, this variant of KE-Opt becomes solvable in polynomial time again, by reducing to a maximum matching problem in a bipartite graph (Abraham et al., , 2007). The problem is also solvable in polynomial time when there are a constant number of vertex types in the compatibility graph (two vertices have the same type if they have the same in- and out-neighbourhoods)  (Dickerson et al., , 2017). However, KE-Opt becomes $\mathcal{NP}$-hard in the case that cycles and chains have unbounded length when recipients and donors can form clubs, such that a club’s donors are only willing to donate kidneys outside of the club if the club’s recipients receive kidneys from outside of the club (Farina et al., , 2017).
@@
Xiao and Wang, (2018) provided an exact algorithm to find a maximum size set of exchanges, with a running time of $O(2^{n}n^{3})$, where $n$ is the number of RDPs. Further approaches to designing exact algorithms are based on parameterised complexity. Lin et al., (2019) developed two randomised fixed-parameter tractable ($\mathcal{FPT}$) algorithms for KE-Opt without chains; the first is for cycles of length at most $3$ and is parameterised by the number of $2$-cycles and $3$-cycles, whilst the second is for cycles of length at most $K$ (for fixed $K\geq 3$) and is parameterised by the number of transplants. Maiti and Dey, (2022) described $\mathcal{FPT}$ algorithms when parameterised by the number of transplants, vertex types, or the sum of the graph’s treewidth and the maximum cycle or chain limit. Hébert-Johnson et al., (2024) showed that the problem is $\mathcal{W}[1]$-hard when parameterised by only treewidth but $\mathcal{FPT}$ when parameterised by vertex type (as defined above). They also gave an improved $\mathcal{FPT}$ algorithm when the parameter is the number of transplants.
@@
Another approach is approximation algorithms. Biró et al., (2009) showed that the problem of finding a maximum size set of exchanges with cycles of length at most $3$ is $\mathcal{APX}$-complete, i.e., an optimal solution cannot be approximated within some constant factor in polynomial time unless $\mathcal{P}=\mathcal{NP}$. Luo et al., (2016) established inapproximability results for KE-Opt in both the weighted (i.e., we seek a maximum weight set of exchanges in the presence of arc weights in the compatibility graph) and unweighted cases, giving specific lower bounds beyond which KE-Opt is not approximable unless $\mathcal{P}=\mathcal{NP}$. Jia et al., (2017) provided a black-box reduction linking the cycle packing problem (a reformulation of the problem of finding a maximum size set of exchanges in a KEP instance) to set packing, leading to a $(3/2+\varepsilon)$-approximation algorithm for cycle limit $3$ and a $(7/3+\varepsilon)$-approximation when there is a cycle length limit of $4$.
@@ 2.2 Hierarchical Optimisation in KE-Opt
For many countries, their KEPs do not solely aim to maximise the number of transplants found in a given matching run. Instead, they typically find a set of exchanges that is optimal with respect to several objectives that take into account the number of transplants selected, the solution’s structure, and notions of fairness and recourse. OR provides a range of methods to optimise over multiple objectives, and many KEPs use hierarchical optimisation (Biró et al., , 2021). This type of multi-objective optimisation sequentially optimises the objectives according to some priority order; when optimising a given objective, the optimal values for the previous objectives are maintained.111 Evidence suggests that all hierarchical objectives can be discriminating in practice. Unpublished simulations carried out by the fifth author on historical data from the UK KEP, which uses five hierarchical objectives, showed that most instances still had over 10,000 feasible solutions after optimising on the first four objectives.
@@
The hierarchical objectives used in KEPs and their orderings can differ greatly between countries; the survey by Biró et al., (2021) highlighted these differences across Europe. For example, the Spanish national KEP uses four objectives and optimises them hierarchically. Given the pool of RDPs, they first find the maximum number of transplants, say $n_{T}$. They then find the maximum number $n_{E}$ of distinct exchanges in a solution with $n_{T}$ transplants (forcing selected chains and cycles to be shorter). Following this, they find a solution with $n_{T}$ transplants, $n_{E}$ distinct exchanges, and a maximum number of cross-arcs. [^2] A cross-arc is an additional arc among the NDD/RDPs in an exchange $c$, but not already an arc of $c$, that allows an alternative exchange $c^{\prime}$ to take place in the case of a vertex or arc failure in $c$ (e.g., if a donor becomes ill, or if a laboratory crossmatch is identified). Including an objective that maximises the number of cross-arcs in a solution gives additional recourse possibilities if some transplants cannot proceed. Finally, while maintaining all the previous optimal values, the chosen solution has an optimal total weight with respect to weights associated with the selected arcs. Kratz, (2024) studied the conditions under which prioritisation of certain types of recipients does not impact the number of transplants identified.
@@
Given that hierarchical optimisation often leads to multiple $\mathcal{NP}$-hard problems being solved in succession, various OR techniques have been employed to find optimal solutions. Manlove and O’Malley, (2014) extended the classical cycle formulation given by Roth et al., (2007) to be able to find solutions according to the UK’s optimality criteria. They then analysed the impact of extending the current criteria to allow cycles of length $4$ instead of $3$ and showed that this could increase the total number of transplants on real data from the UK’s national KEP. Glorie et al., 2014b , created an iterative branch-and-price algorithm to find optimal exchanges for the Dutch national KEP criteria (see also Section 3.3). Klimentova et al., (2014) also studied a hierarchical optimisation problem that involves first maximising the number of transplants and then maximising the number of exchanges (a proxy for minimising the average exchange length). They solved this problem by creating a single combined objective function incorporating a suitable choice of weights , so that the individual objective functions are optimised hierarchically, using a branch-and-price algorithm. Interestingly, they obtained faster solution times by optimising on the single combined objective, compared to solving sequentially on the individual objectives. Delorme et al., (2024) developed four strategies to reduce the running time of ILP-based algorithms when dealing with hierarchical optimisation: eliminating the dominated exchanges (cycles or chains that cannot appear in any optimal solution given the considered criteria), objective diving (using dual bounds to set the value of early objectives, possibly backtracking if a fixed value is, in fact, infeasible), reduced cost variable fixing (see Section 3.3), and model swapping (using a different ILP model depending on the objective currently optimised). The effectiveness of their approach was demonstrated on three sets of hierarchical objective functions: from the UK, Spain, and the Netherlands.
@@ 2.3 Enabling Equal Access to Transplantation
The next area of the literature we explore relates to ensuring that recipients have equal access to transplantation, and thus, many of the cited works provide an economic perspective on KEPs. There can be many reasons why some recipients could be less likely to be selected for a transplant in a KEP. So-called hard-to-match recipients tend to have fewer compatible donors in the pool, and when maximising the number of transplants in a matching run, these recipients could repeatedly not be selected, raising concerns about fairness.
@@
One way to enable equal access to transplantation involves altering the mechanism to prioritise the likelihood of certain recipients being selected. Roth et al., (2005) were the first to study algorithmic fairness in KEPs. Their egalitarian mechanism maximises the potential utility of harder-to-match RDPs when KEPs allow for stochastic mechanisms, i.e. , there is a probability distribution over feasible sets of exchanges. Li et al., (2014) provided polynomial time algorithms to combat the exponential algorithm given by Roth et al., (2005), while keeping a level of equity among the RDPs in their likelihood of being selected.
Another aspect of fairness relates to the likelihood of receiving a transplant when there are multiple optimal solutions, and one must be selected. Farnadi et al., (2021) focused on the equity of access to transplantation, especially in the presence of multiple optimal solutions. They enumerated all optimal solutions by using a hybrid of constraint programming and linear programming to avoid any bias that a deterministic solver may have. To efficiently compute a fair lottery over multiple feasible sets of exchanges, St-Arnaud et al., (2025) proposed mixed-integer conic programs modelling fair lotteries and described the use of column generation to solve them. Demeulemeester et al., (2025) also studied fairness in KEPs when there are multiple optimal solutions and the solution is selected by a general-purpose mathematical solver. They demonstrated that not all optimal solutions had the same probability of being returned by the solver and proposed several algorithms to alleviate this issue. They also extended the discussion of multiple optimal solutions to general ILP problems.
@@
Highly sensitised recipients have a low level of tissue-type compatibility and therefore, tend to be harder to match. As such, studies have focused on mechanisms that relax the objective to maximise the number of transplants, in order to help these recipients. Dickerson et al., (2014) were the first to introduce the price of fairness (PoF) to KEPs, a measure that quantifies the trade-off between efficiency (maximising the number of transplants) and fairness (improving access to transplantation for highly-sensitised recipients). McElfresh and Dickerson, (2018) extended upon this by studying the PoF when there are also NDDs, showing that the PoF tends to zero when there are many NDDs. On the other hand, Ashlagi et al., (2012) showed through theoretical results and via simulations that longer chains could help select more highly sensitised recipients without negatively impacting the likelihood of the other RDPs from being selected. However, McElfresh and Dickerson, (2018) showed that the mechanisms by Dickerson et al., (2014) may have an arbitrarily bad PoF when the length of cycles and chains increase. Yet, their findings suggest that fairer solutions for highly sensitised patients tend to use longer chains and cycles. The mechanisms created by McElfresh and Dickerson, (2018) were shown to limit the loss in efficiency in terms of PoF when directly prioritising disadvantaged recipients. Duppala et al., (2023) presented randomised polynomial-time algorithms that take into account proportionality vectors that indicate how to prioritise certain groups, returning a probabilistically fair solution with provable guarantees.